OpenStudy (thomasfire55):

Use the rules of exponents to evaluate or simplify. Write without negative exponents. (3)4^0

OpenStudy (thomasfire55):

@TheSmartOne

TheSmartOne (thesmartone):

is that \(\Large (3)4^0\)

OpenStudy (thomasfire55):

yes

jhonyy9 (jhonyy9):

so just you need to know that every numbers on exponent zero is equal 1

jhonyy9 (jhonyy9):

so 4^0 = 1

OpenStudy (thomasfire55):

so the awnser is 3?

jhonyy9 (jhonyy9):

than there are 3*4^0 = 3*1 = 3 yes exactly

OpenStudy (thomasfire55):

thanks

jhonyy9 (jhonyy9):

my pleasure good luck